# Asymptotic stability equals exponential stability—while you twist your eyes

### v1

@@ 3   Stability preserving homotopies
Note that $\widetilde{\varphi}^{t}(\cdot;0)=\varphi^{t}(\cdot;-\nabla V)$ whereas $\widetilde{\varphi}^{t}(\cdot;1)=T\circ\varphi^{t}(\cdot;-\nabla V)\circ T^{-1}$. Also note that continuity of $[0,1]\ni s\mapsto\widetilde{\varphi}^{t}(\cdot;s)$, in particular, continuity of $[0,1]\ni s\mapsto\widetilde{H}(s,\cdot)^{-1}$, follows from $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$ being a topological group [ref:arens1946topologies] (endowed with the compact-open topology), i.e., combine that the map $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})\ni h\mapsto h^{-1}$ is continuous with path-connectedness being preserved under a continuous map. Then, it follows from (3.1) that for any $t>0$ and $x\neq 0$ we have that
$V(\widetilde{H}(s,\widetilde{\varphi}^{t}(x;s)))$ $-V(\widetilde{H}(s,\widetilde{\varphi}^{0}(x;s)))\leq$ (3.2)
$-\int^{t}_{0}W_{V}(\widetilde{H}(s,\widetilde{\varphi}^{\tau}(x;s)))\mathrm{d}\tau<0.$
@@
Now, using (3.2), define the (parametric in $s\in[0,1]$) Lyapunov function $\widetilde{V}(\cdot;s)\in C^{0}(\mathbb{R}^{n};\mathbb{R}_{\geq 0})$ through $x\mapsto\widetilde{V}(x;s):=V(\widetilde{H}(s,x))$ and similarly, define $\widetilde{W}_{V}$ through $x\mapsto\widetilde{W}_{V}(x;s):=W_{V}(\widetilde{H}(s,x))$ Hence, we have that for any $s\in[0,1]$ the following Lyapunov inequality holds: $\widetilde{V}(\widetilde{\varphi}^{t}(x;s);s)-\widetilde{V}(\widetilde{\varphi}^{0}(x;s);s)\leq-\int^{t}_{0}\widetilde{W}_{V}(\widetilde{\varphi}^{\tau}(x;s);s)\mathrm{d}\tau<0$, for all $t>0$ and any $x\neq 0$. Note that $x\mapsto\widetilde{V}(x;0)=V(x)$ whereas $x\mapsto\widetilde{V}(x;1)=V(T^{-1}(x))=\gamma(\|x\|_{2})=:V_{\gamma}(x)$ and $x\mapsto\widetilde{W}_{V}(x;1)=W_{V}(T^{-1}(x))$. Also, as $\widetilde{H}(s,\cdot)\in\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$ fixes 00, any compact neighbourhood $K$ of 00 is mapped to some compact neighbourhood $\widetilde{H}(s,K)$ of 00, hence, $\widetilde{V}(\cdot;s)$ will always have the required compact sublevel sets.
@@
(iv) Although the path from $\mathrm{id}_{\mathbb{R}^{n}}$ to $T^{-1}$ is merely through $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$, it is known [^7] Some comments are in place. Perelman’s resolution came after the paper by Grüne, Sontag and Wirth, the case $n=4$ ($\mathbb{S}^{3}$) is resolved by now. Despite some claims in the literature, to the best of our knowledge, the case $n=5$ is still open. We emphasize that the existence of these diffeomorphisms is largely due to Smale’s h-cobordism theorem, as the generalized Poincaré conjecture is known to fail, in general, for $\mathsf{Diff}$ (e.g., due to the existence of exotic spheres [ref:Milnor7]). that $T$ can always be chosen to be diffeomorphic on $\mathbb{R}^{n}\setminus\{0\}$ for $n\neq 5$ [ref:grune1999asymptotic] (note that the composition with $\rho$ does not change this). Hence, the vector field corresponding to $\widetilde{\varphi}(\cdot;1)$, denoted $\widetilde{F}$, is readily well-defined on $\mathbb{R}^{n}\setminus\{0\}$, i.e.,
@@
In general, $\widetilde{F}$ need not be continuous at 00, simply because $\mathrm{D}T$ need not be continuous at 00. However, $\gamma$ can always be chosen such that $T$ is $C^{1}$ on $\mathbb{R}^{n}$ with $\mathrm{D}T(0)=0$ [ref:grune1999asymptotic, Prop. 1].
Unfortunately, $V_{\gamma}$ is by no means smooth at 00 for any choice of $\gamma$. In fact, the appropriate $\gamma\in\mathcal{K}_{\infty}$ from [ref:grune1999asymptotic, Prop. 1] to guarantee $T$ is sufficiently regular is of the following form. Let $\alpha\in\mathcal{K}$ be smooth, define $h(r):=\int^{r}_{0}a(\tau)\mathrm{d}\tau$ and set $\gamma:=h^{-1}$. It readily follows that $\gamma^{\prime}(r)=1/\alpha(\gamma(r))$ is smooth on $(0,+\infty)$, yet, $\lim_{r\to 0^{+}}\gamma^{\prime}(r)=+\infty$, e.g., $r\mapsto\gamma(r)=r^{1/2}$. However, note that on $\mathbb{R}^{n}\setminus\{0\}$ we have $\langle\nabla V_{\gamma}(x),\widetilde{F}(x)\rangle\leq-\widetilde{W}_{V}(x)$, which is equivalent to
@@
Thus, multiplying (3.4) by the function $a\in C^{0}(\mathbb{R}^{n};\mathbb{R}_{\geq 0})$, defined through $x\mapsto a(x):=\|x\|_{2}/\gamma^{\prime}(\|x\|_{2})$, yields that $V_{q}\in C^{\infty}(\mathbb{R}^{n};\mathbb{R}_{\geq 0})$ and $W_{q}\in C^{\infty}(\mathbb{R}^{n}\setminus\{0\};\mathbb{R}_{>0})$, defined by $x\mapsto V_{q}(x):=\frac{1}{2}\|x\|_{2}^{2}$ and $x\mapsto W_{q}(x):=-a(x)\widetilde{W}_{V}(x)$, comprise a Lyapunov pair for $\widetilde{F}$, also recall Example 1.4. Equivalently, one can observe that the level sets of $V_{\gamma}$ are standard spheres.
Then, as in Step (i), we construct a homotopy through non-vanishing vector fields on $\mathbb{R}^{n}\setminus\{0\}$, in this case from $\widetilde{F}$ to $-\nabla V_{q}=-\partial_{x}$, with $V_{q}$ being a Lyapunov function asserting that 00 is (strongly) GAS under $\widetilde{F}$. Specifically, construct the map $(s,x)\mapsto H_{q}(s,x):=(1-s)\widetilde{F}(x)-s\nabla V_{q}(x)$ (in this case for all $x\in\mathbb{R}^{n}$). Then, as for any $s\in[0,1]$ we have that $(1-s)\widetilde{F}-s\nabla V_{q}\in C^{0}(\mathbb{R}^{n};\mathbb{R}^{n})$, plus
$\langle\nabla V_{q}(x),H_{q}(s,x)\rangle\leq-(1-s){W}_{q}(x)-2sV_{q}(x),\,\,\forall\,x\in\mathbb{R}^{n}\setminus\{0\},$
we find that 00 is (strongly) GAS under $\dot{x}=(1-s)\widetilde{F}(x)-s\nabla V_{q}(x)$ for any fixed $s\in[0,1]$. To pass to the level of semiflows we can proceed as in Step (i), that is, we can show there is homotopy from $\varphi^{t}(\cdot;\widetilde{F})$ to $\varphi^{t}(\cdot;-\partial_{x})$ along semiflows, such that 00 is GAS.
@@
Proposition 3.2 is largely about the existence of a homotopy, less about the construction. Still, we provide some examples below.
@@
It is known that the following polynomial dynamical system does not admit a polynomial Lyapunov function asserting 00 is GAS [ref:ahmadi2011globally]. This $2$-dimensional differential equation is defined by $\dot{x}=X_{1}(x):=(-x_{1}+x_{1}x_{2},-x_{2})$. A non-polynomial $C^{\infty}$ Lyapunov function that does assert 00 is GAS is given by $(x_{1},x_{2})\mapsto V_{1}(x_{1},x_{2}):=(1/2)\log(1+x_{1}^{2})+(1/2)x_{2}^{2}$. To construct a homotopy from $\varphi(\cdot;X_{1})$ to $\varphi(\cdot;-\partial_{x})$, first construct the straight-line homotopy from $X_{1}$ to $-\nabla V_{1}$, i.e. $(1-s)X_{1}-s\nabla V_{1}$; stability is asserted throughout by $V_{1}$. Next, define $V(\cdot;s):=(1-s)V_{1}+sV_{q}$, for $s\in[0,1]$ and $x\mapsto V_{q}(x):=(1/2)\|x\|_{2}^{2}$. It follows that, for any $s\in[0,1]$, $\nabla V(x;s)=0\iff x=0$, such that $[0,1]\ni s\mapsto-\nabla V(\cdot;s)$ parametrizes a homotopy from $-\nabla V_{1}$ to $-\partial_{x}$, along vector fields that render 00 GAS, asserted by $V(\cdot;s)$. It is interesting to note that although $X_{1}$ does not admit a convex polynomial Lyapunov function, the negative gradient flow of the Lyapunov function corresponding to $X_{1}$ does cf[ref:JongeneelSchwan2024TAC]. $\circ$
Figure 3.1: Example 3.2: on the left, the graph of $V_{i}$ around $x=0$; and on the right, the graph of $V_{q}$ around $x=0$. In between, two steps of the homotopy that connects $V_{i}$ to $V_{q}$, through continuous function that preserve $x=0$ being the global minimizer.
@@
Consider a coercive, invex function (i.e., every critical point is a global minimizer) on $\mathbb{R}$ defined by $x\mapsto V_{i}(x):=(1/2)x^{2}+(3/2)\sin(x)^{2}$. In the context of Proposition 3.2, we can find a map $T$ such that $V(T^{-1}(x))=\gamma(|x|):=\sqrt{|x|}$, and a path from $T$ to $\mathrm{id}_{\mathbb{R}}$, as given by the homotopy
$[0,1]\times\mathbb{R}\ni(s,x)\mapsto H_{i}(s,x):=\mathrm{sgn}(x)\left(\frac{1+s}{2}x^{2}+(1-s)\frac{3}{2}\sin(x)^{2}\right)^{2-(3/2)s}$
with $H_{i}(0,\cdot)=T$ and $H_{i}(1,\cdot)=\mathrm{id}_{\mathbb{R}}$. In a second step one could consider the homotopy $(s,x)\mapsto x^{(1/2)+(3/2)s}$. Summarizing, we can construct the path from $V_{i}$ to $x\mapsto V_{q}(x):=x^{2}$, along continuous functions such that 00 is the global minimizer throughout [^8] For a simulation of this homotopy, see wjongeneel.nl/figinvex.gif., see Figure 3.1. Evidently, for examples as simple as $V_{i}$ one can find simpler homotopies. However, to the best of our knowledge, the mere existence of such a homotopy is still an open problem, in general.
@@
However, Proposition 3.2 provides us with the existence of a homotopy from a smooth, proper Lyapunov function $V$ on $\mathbb{R}^{n}$, asserting $0\in\mathbb{R}^{n}$ is GAS, to $\mathbb{R}^{n}\ni x\mapsto\|x\|_{2}^{2}$, along continuous [^9] It is not evident, and currently unknown, whether smoothness can be preserved throughout the homotopy, see Step (iii) of the proof of Proposition 3.2. Lyapunov functions that assert $0\in\mathbb{R}^{n}$ is GAS along the homotopy. Differently put, we can find a homotopy from a coercive, invex function, to a convex function, such that along the homotopy we preserve the global minimizer. We believe this is of independent interest. $\circ$
@@
As we allow for a class of discontinuous vector fields, Proposition 3.2 allows for a slight extension of the standard Hopf index, this has been pioneered by Gottlieb in earlier work, e.g., see [ref:gottlieb1995index]. See also [ref:casagrande2008conley] for a Conley index applicable to discontinuous vector fields and see [ref:Kvalheim2021PHhybrid] for a hybrid version of the Poincaré-Hopf theorem.
We point out that one could omit Step (i) of the proof of Proposition 3.2, yet, as is also evident from Example 3.1, it is typically convenient to pass through a gradient flow. We exploit precisely this step in the proof of Theorem 3.4 below.
@@
We also remark that the origin plays no particular role in Proposition 3.2, as it should. In fact, if $X$ is such as in Proposition 3.2, yet, the equilibrium point is now arbitrary, we can still construct a homotopy between $\varphi(\cdot;\mathcal{F}[X])$ and $\varphi(\cdot;-\partial_{x})$ such that along the homotopy some point is GAS. For instance, the following family of time-$t$ maps parametrizes a homotopy between the flows corresponding to the ODEs $\dot{x}=-x$ and $\dot{x}=-(x-\bar{x})$:
$(t,x)\mapsto\varphi^{t}(x;s)=e^{-t}x+s(I_{n}-e^{-t}I_{n})\bar{x},\quad s\in[0,1].$
Indeed, $s\bar{x}$ is GAS along the homotopy (e.g., consider the Lyapunov function $x\mapsto V(x;s):=\frac{1}{2}(x-s\bar{x})^{2}$).
@@
Next, we generalize Proposition 3.2 to smooth manifolds, which is almost immediate as points being GAS heavily restrict the class of manifolds, e.g., see [ref:bhatia1970stability, Ch. V.3]. To clarify terminology, an $n$-dimensional manifold ${M}^{n}$ is said to be $\psi$-diffeomorphic to $\mathbb{R}^{n}$ when $\psi({M}^{n})=\mathbb{R}^{n}$ for a diffeomorphism $\psi$. We will always assume that our manifold is Hausdorff and second countable. Regarding our vector fields, we assume the following.
@@
Given a $C^{\infty}$ manifold $M^{n}$, the continuous vector field $X$ on $M^{n}$ is such that $X$ gives rise to a continuous (global) flow $\varphi(\cdot;X):\mathbb{R}\times M^{n}\to M^{n}$.
Comparing to Proposition 3.2, we note that the vector fields under consideration in Theorem 3.4 (i.e., Assumption 3.3) can be less regular indeed. To avoid too many technicalities, we refrain from those generalizations, however.
@@
Let $M^{n}$ be a $C^{\infty}$ manifold, for $n\neq 5$, and let $p^{\star}\in M^{n}$ be GAS under both the vector fields $X$ and $Y$, with both $X$ and $Y$ satisfying Assumption 3.3. Then, the flow $\varphi(\cdot;X)$ is homotopic to $\varphi(\cdot;Y)$, along flows that preserve $p^{\star}$ being GAS.
@@
First, since $p^{\star}$ is GAS under $X$, there is $V\in C^{\infty}(M^{n};\mathbb{R}_{\geq 0})$ such that $X(V)=L_{X}V<0$ on $M^{n}\setminus\{p^{\star}\}$ [ref:fathi2019smoothing]. Now we fix some Riemannian metric $g$, which always exists [Lee2, Ch. 13], such that we can define the Riemannian gradient $\mathrm{grad}\,V$ (omitting the dependence on $g$). It follows that $-\mathrm{grad}\,V(V)<0$ on $M^{n}\setminus\{p^{\star}\}$ such that we can define the straight-line homotopy between $X$ and $-\mathrm{grad}\,V$, preserving that $p^{\star}$ is GAS.
Due to smoothness and stability, we know that $-\mathrm{grad}\,V$ gives rise to a flow, denoted $\varphi(\cdot;-\mathrm{grad}\,V)$, on $M^{n}$. In fact, for any $t\in\mathbb{R}$, $p\mapsto\varphi^{t}(p;-\mathrm{grad}\,V)$ is a diffeomorphism. This implies in particular that $M^{n}$ is $\psi$-diffeomorphic to $\mathbb{R}^{n}$ [ref:wilson1967structure, Thm. 2.2], for some diffeomorphism $\psi:M^{n}\to\mathbb{R}^{n}$.
Next, define the diffeomorphism $\widetilde{\psi}$ by ${M}^{n}\ni p\mapsto\widetilde{\psi}(p):=\psi(p)-\psi(p^{\star})\in\mathbb{R}^{n}$, such that $\widetilde{\psi}(p^{\star})=0$, and consider $x:=\widetilde{\psi}(p)$ such that in these new coordinates we have
$\dot{x}=-\mathrm{D}\widetilde{\psi}(\widetilde{\psi}^{-1}(x))\mathrm{grad}\,V(\widetilde{\psi}^{-1}(x))=:X(x)$
Indeed, the vector field $X$ meets precisely the criteria of Proposition 3.2. Then, observe that $\varphi^{t}(\cdot;-\mathrm{grad}\,V)=\widetilde{\psi}^{-1}\circ\varphi^{t}(\cdot;X)\circ\widetilde{\psi}$, which implies that $\varphi^{t}(\cdot;-\mathrm{grad}\,V)$ can be homotoped to $\widetilde{\psi}^{-1}\circ\varphi^{t}(\cdot;-\partial_{x})\circ\widetilde{\psi}$ such that $p^{\star}$ remains GAS along the homotopy.
We can do the exact same for the vector field $Y$. Then, since $\widetilde{\psi}$ is independent of the vector fields $X$ and $Y$ we can conclude by the transitivity property of homotopies, that is, both the time-$t$ maps under $X$ and $Y$ can be homotoped to $\widetilde{\psi}^{-1}\circ\varphi^{t}(\cdot;-\partial_{x})\circ\widetilde{\psi}$.
@@
Example 3.3 (Equilibria on the sphere).
Regarding examples of Theorem 3.4 one might think of rendering the South Pole $S$ an attractor on $\mathbb{S}^{2}\setminus\{N\}$. To that end, consider the vector fields $X$ and $Y$ on $\mathbb{R}^{2}$, defined through

### v2

@@ 3   Stability preserving homotopies
Note that $\widetilde{\varphi}^{t}(\cdot;0)=\varphi^{t}(\cdot;-\nabla V)$ whereas $\widetilde{\varphi}^{t}(\cdot;1)=T\circ\varphi^{t}(\cdot;-\nabla V)\circ T^{-1}$. Also note that continuity of $[0,1]\ni s\mapsto\widetilde{\varphi}^{t}(\cdot;s)$, in particular, continuity of $[0,1]\ni s\mapsto\widetilde{H}(s,\cdot)^{-1}$, follows from $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$ being a topological group [ref:arens1946topologies] (endowed with the compact-open topology), i.e., combine that the map $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})\ni h\mapsto h^{-1}$ is continuous with path-connectedness being preserved under a continuous map. Then, by (3.1), we get
$V(\widetilde{H}(s,\widetilde{\varphi}^{t}(x;s)))-V(\widetilde{H}(s,\widetilde{\varphi}^{0}(x;s)))\leq-\int^{t}_{0}W_{V}(\widetilde{H}(s,\widetilde{\varphi}^{\tau}(x;s)))\mathrm{d}\tau.$ (3.2)
@@
Now, using (3.2), define $\widetilde{V}(\cdot;s)\in C^{0}(\mathbb{R}^{n};\mathbb{R}_{\geq 0})$ through $x\mapsto\widetilde{V}(x;s):=V(\widetilde{H}(s,x))$ and similarly, define $x\mapsto\widetilde{W}_{V}(x;s):=W_{V}(\widetilde{H}(s,x))$ Hence, for any $s\in[0,1]$, we have $\widetilde{V}(\widetilde{\varphi}^{t}(x;s);s)-\widetilde{V}(\widetilde{\varphi}^{0}(x;s);s)\leq-\int^{t}_{0}\widetilde{W}_{V}(\widetilde{\varphi}^{\tau}(x;s);s)\mathrm{d}\tau<0$, $\forall(t,x)\in\mathbb{R}_{>0}\times\mathbb{R}^{n}\setminus\{0\}$. Note that $x\mapsto\widetilde{V}(x;0)=V(x)$, $x\mapsto\widetilde{V}(x;1)=V(T^{-1}(x))=\gamma(\|x\|_{2})=:V_{\gamma}(x)$ and $x\mapsto\widetilde{W}_{V}(x;1)=W_{V}(T^{-1}(x))$. Also, as $\widetilde{H}(s,\cdot)$ fixes 00, $\widetilde{V}(\cdot;s)$ will always have the required compact sublevel sets.
@@
(iv) Although the path from $\mathrm{id}_{\mathbb{R}^{n}}$ to $T^{-1}$ is merely through $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$, it is known that $T$ can be chosen to be diffeomorphic on $\mathbb{R}^{n}\setminus\{0\}$ for $n\neq 5$ [ref:grune1999asymptotic] (the composition with $\rho$ does not change this) [^6] Some comments are in place. Perelman’s work came after [ref:grune1999asymptotic], resolving the case $n=4$ ($\mathbb{S}^{3}$). Despite some claims in the literature, to the best of our knowledge, the case $n=5$ ($\mathbb{S}^{4}$) is still open. As the generalized Poincaré conjecture is known to fail, in general, for $\mathsf{Diff}$ (e.g., due to the existence of exotic spheres [ref:Milnor7]), we emphasize that the existence of these diffeomorphisms is due to Smale’s h-cobordism theorem (e.g., see [ref:smale1962structure, Thm. 5.1]). To be more explicit, we can construct a h-cobordism between $V^{-1}(c)$ and a levelset of $x\mapsto\|x\|_{2}^{2}$, e.g., by appealing to the same type of arguments as in [ref:jongeneelECC24, Lem. III.4].. Hence, the vector field corresponding to $\widetilde{\varphi}(\cdot;1)$, denoted $\widetilde{F}$, is well-defined on $\mathbb{R}^{n}\setminus\{0\}$, i.e.,
@@
In general, $\widetilde{F}$ need not be continuous at 00, this, because $\mathrm{D}T$ need not be continuous at 00. However, $\gamma$ can always be chosen such that $T$ is $C^{1}$ on $\mathbb{R}^{n}$ with $\mathrm{D}T(0)=0$ [ref:grune1999asymptotic, Prop. 1]. Unfortunately, $V_{\gamma}$ is also by no means smooth at 00 for any choice of $\gamma$. In fact, the appropriate $\gamma\in\mathcal{K}_{\infty}$ to guarantee that $T$ is sufficiently regular is of the following form. Let $\alpha\in\mathcal{K}$ be smooth, define $h(r):=\int^{r}_{0}a(\tau)\mathrm{d}\tau$ and set $\gamma:=h^{-1}$. It follows that $\gamma^{\prime}(r)=1/\alpha(\gamma(r))$ is smooth on $(0,+\infty)$, yet, $\lim_{r\to 0^{+}}\gamma^{\prime}(r)=+\infty$, e.g., $r\mapsto\gamma(r)=r^{1/2}$. However, see that on $\mathbb{R}^{n}\setminus\{0\}$ we have $\langle\nabla V_{\gamma}(x),\widetilde{F}(x)\rangle\leq-\widetilde{W}_{V}(x)$, that is,
@@
Thus, multiplying (3.4) by the function $a\in C^{0}(\mathbb{R}^{n};\mathbb{R}_{\geq 0})$, defined through $x\mapsto a(x):=\|x\|_{2}/\gamma^{\prime}(\|x\|_{2})$, yields that $V_{q}\in C^{\infty}(\mathbb{R}^{n};\mathbb{R}_{\geq 0})$ and $W_{q}\in C^{\infty}(\mathbb{R}^{n}\setminus\{0\};\mathbb{R}_{>0})$, defined by $x\mapsto V_{q}(x):=\frac{1}{2}\|x\|_{2}^{2}$ and $x\mapsto W_{q}(x):=-a(x)\widetilde{W}_{V}(x)$, comprise a $C^{\infty}$ Lyapunov pair for $\widetilde{F}$, also recall Example 1.5. Equivalently, one can observe that the level sets of $V_{\gamma}$ are standard spheres.
Then, as in Step (i), construct a homotopy through non-vanishing vector fields on $\mathbb{R}^{n}\setminus\{0\}$, in this case from $\widetilde{F}$ to $-\nabla V_{q}=-\partial_{x}$, with $V_{q}$ being a Lyapunov function asserting stability of 00 throughout. Then, we show there is homotopy from $\varphi^{t}(\cdot;\widetilde{F})$ to $\varphi^{t}(\cdot;-\partial_{x})$ along semiflows, such that 00 is GAS.
@@
As Proposition 3.2 is about existence, we provide some examples below.
@@
It is known that the polynomial ODE $\dot{x}=X_{1}(x):=(-x_{1}+x_{1}x_{2},-x_{2})$ does not admit a polynomial Lyapunov function asserting 00 is GAS [ref:ahmadi2011globally]. A non-polynomial Lyapunov function, however, is given by $(x_{1},x_{2})\mapsto V_{1}(x_{1},x_{2}):=\tfrac{1}{2}\log(1+x_{1}^{2})+\tfrac{1}{2}x_{2}^{2}$. To construct a homotopy from $\varphi(\cdot;X_{1})$ to $\varphi(\cdot;-\partial_{x})$, first construct the straight-line homotopy from $X_{1}$ to $-\nabla V_{1}$, with stability asserted throughout by $V_{1}$. Next, define $V(\cdot;s):=(1-s)V_{1}+sV_{q}$, for $s\in[0,1]$ and $x\mapsto V_{q}(x):=\tfrac{1}{2}\|x\|_{2}^{2}$. As for any $s\in[0,1]$, $\nabla V(x;s)=0\iff x=0$, $[0,1]\ni s\mapsto-\nabla V(\cdot;s)$ parametrizes a homotopy from $-\nabla V_{1}$ to $-\partial_{x}$, along vector fields that render 00 GAS, asserted by $V(\cdot;s)$. Although $X_{1}$ does not admit a convex polynomial Lyapunov function, the negative gradient flow of the Lyapunov function corresponding to $X_{1}$ does cf[ref:JongeneelSchwan2024TAC]. $\circ$
Figure 3.1: Example 3.4: on the left, the graph of $v_{i}$ around $x=0$; and on the right, the graph of $v_{q}$ around $x=0$. In between, steps of the homotopy that connects $v_{i}$ to $v_{q}$, while preserving that $x=0$ is the global minimizer.
The purpose of the following example is to illustrate the construction of the homotopy that appears in the proof of Proposition 3.2.
@@
Consider a coercive, invex function (i.e., every critical point is a global minimizer) on $\mathbb{R}$ defined by $x\mapsto v_{i}(x):=\tfrac{1}{2}x^{2}+\tfrac{3}{2}\sin(x)^{2}$. Along the lines of the proof of Proposition 3.2, we can find a map $T$ such that $v_{i}(T^{-1}(x))=\gamma(|x|):=\sqrt{|x|}$ (see that $\gamma(s)/\gamma^{\prime}(s)\geq s$ for all $s\geq 0$), and a path from $T$ to $\mathrm{id}_{\mathbb{R}}$, as given by
$[0,1]\times\mathbb{R}\ni(s,x)\mapsto H_{i}(s,x):=\mathrm{sgn}(x)\left(\frac{1+s}{2}x^{2}+(1-s)\frac{3}{2}\sin(x)^{2}\right)^{2-(3/2)s},$
@@
with $H_{i}(0,\cdot)=T$ and $H_{i}(1,\cdot)=\mathrm{id}_{\mathbb{R}}$ (e.g., $v_{i}=\sqrt{|H_{i}(0,\cdot)|}$). Next, consider the homotopy $(s^{\prime},x)\mapsto x^{1/2+(3/2)s^{\prime}}$ to construct the path from $v_{i}$ to $x\mapsto v_{q}(x):=x^{2}$, along continuous functions such that 00 is the global minimizer throughout [^7] For a simulation of this homotopy, see wjongeneel.nl/figinvex.gif., see Figure 3.1. For functions as simple as $v_{i}$ one can find simpler homotopies (e.g., a straight-line), however, to the best of our knowledge, being able to guarantee the mere existence of such a homotopy is new. Proposition 3.2 provides us with the existence of a homotopy from a smooth Lyapunov function $V$ to $V_{q}$, along continuous [^8] It is not evident, and currently unknown, whether smoothness can be preserved throughout the homotopy, see Step (iii) of the proof of Proposition 3.2. Lyapunov functions that assert $0\in\mathbb{R}^{n}$ is GAS along the homotopy. Differently put, we can find a homotopy from a coercive, invex function, to a convex function, such that along the homotopy the minimizer is preserved. This might be of independent interest. $\circ$
@@
As we allow for a class of discontinuous vector fields, Proposition 3.2 allows for an extension of the standard Hopf index, this has been pioneered by Gottlieb, e.g., see [ref:gottlieb1995index]. See also [ref:casagrande2008conley] for a Conley index applicable to discontinuous vector fields and see [ref:Kvalheim2021PHhybrid] for a hybrid Poincaré-Hopf theorem.
@@
We point out that one could omit Step (i) of the proof of Proposition 3.2, yet, as is also shown in Example 3.3, it is typically convenient to pass through a gradient flow. We also remark that the origin plays no particular role in Proposition 3.2, as it should. In fact, if $X$ is such as in Proposition 3.2, yet, the equilibrium point is now arbitrary, we can still construct a homotopy between $\varphi(\cdot;\mathcal{F}[X])$ and $\varphi(\cdot;-\partial_{x})$ such that along the homotopy some point is GAS. For instance, the following family of time-$t$ maps parametrizes a homotopy between the flows corresponding to the ODEs $\dot{x}=-x$ and $\dot{x}=-(x-\bar{x})$: $(t,x)\mapsto\varphi^{t}(x;s)=e^{-t}x+s(I_{n}-e^{-t}I_{n})\bar{x}$, for $s\in[0,1]$. Indeed, $s\bar{x}$ is GAS along the homotopy (e.g., consider the Lyapunov function $x\mapsto V(x;s):=\frac{1}{2}(x-s\bar{x})^{2}$).
@@
Next, we generalize Proposition 3.2 to smooth manifolds, which is almost immediate as the GAS property heavily restricts the class of manifolds, e.g., see [ref:bhatia1970stability, Ch. V.3]. We assume that our manifolds are Hausdorff and second countable. Regarding our vector fields, we assume simply the following.
@@
Our vector fields are locally Lipschitz on $M^{n}$.
Indeed, one can work with less regular vector fields and fully generalize Proposition 3.2 to manifolds, i.e., via [ref:mayhew2011topological, Cor. 13]. We refrain from introducing further technicalities and work with Assumption 3.5.
@@
Let $M^{n}$ be a smooth manifold, for $n\neq 5$, and let $p^{\star}\in M^{n}$ be GAS under both the vector fields $X$ and $Y$, with both $X$ and $Y$ satisfying Assumption 3.5. Then, the flow $\varphi(\cdot;X)$ is homotopic to $\varphi(\cdot;Y)$, along semiflows that preserve $p^{\star}$ being GAS.
@@
First, since $p^{\star}$ is GAS under $X$, there is $V\in C^{\infty}(M^{n};\mathbb{R}_{\geq 0})$ such that $X(V)=L_{X}V<0$ on $M^{n}\setminus\{p^{\star}\}$ [ref:fathi2019smoothing]. Now fix some Riemannian metric, which exists [Lee2, Ch. 13], and define the Riemannian gradient $\mathrm{grad}\,V$. It follows that $-\mathrm{grad}\,V(V)<0$ on $M^{n}\setminus\{p^{\star}\}$. Then, for any $t\in\mathbb{R}$, the corresponding time-$t$ map $p\mapsto\varphi^{t}(p;-\mathrm{grad}\,V)$ is a diffeomorphism. This implies there is a diffeomorphism $\psi:M^{n}\to\mathbb{R}^{n}$ [ref:wilson1967structure, Thm. 2.2]. Next, define the diffeomorphism $\widetilde{\psi}$ by ${M}^{n}\ni p\mapsto\widetilde{\psi}(p):=\psi(p)-\psi(p^{\star})\in\mathbb{R}^{n}$, such that $\widetilde{\psi}(p^{\star})=0$. Now set $x:=\widetilde{\psi}(p)$ and consider
$\dot{x}=-\mathrm{D}\widetilde{\psi}(\widetilde{\psi}^{-1}(x))X(\widetilde{\psi}^{-1}(x))=:X_{0}(x).$
Indeed, the vector field $X_{0}$ meets the criteria of Proposition 3.2. Then, observe that $\varphi^{t}(\cdot;X)=\widetilde{\psi}^{-1}\circ\varphi^{t}(\cdot;X_{0})\circ\widetilde{\psi}$, which implies that $\varphi^{t}(\cdot;X)$ can be homotoped to $\widetilde{\psi}^{-1}\circ\varphi^{t}(\cdot;-\partial_{x})\circ\widetilde{\psi}$ such that $p^{\star}$ remains GAS along the homotopy. We can do the exact same for the vector field $Y$. Then, since $\widetilde{\psi}$ works for both $X$ and $Y$ we can conclude by transitivity. ∎
@@
Example 3.7 (Equilibria on $\mathbb{S}^{2}$).
To exemplify Theorem 3.6, one might think of rendering the South Pole $S$ an attractor on $\mathbb{S}^{2}\setminus\{N\}$, for $N$ the North Pole. To that end, consider the vector fields $X$ and $Y$ on $\mathbb{R}^{2}$, defined through